### v1

# Data-Driven Lipschitz Continuity: A Cost-Effective Approach to Improve Adversarial Robustness

Abstract
The security and robustness of deep neural networks (DNNs) have become increasingly concerning. This paper aims to provide both a theoretical foundation and a practical solution to ensure the reliability of DNNs. We explore the concept of Lipschitz continuity to certify the robustness of DNNs against adversarial attacks, which aim to mislead the network with adding imperceptible perturbations into inputs. We propose a novel algorithm that remaps the input domain into a constrained range, reducing the Lipschitz constant and potentially enhancing robustness. Unlike existing adversarially trained models, where robustness is enhanced by introducing additional examples from other datasets or generative models, our method is almost cost-free as it can be integrated with existing models without requiring re-training. Experimental results demonstrate the generalizability of our method, as it can be combined with various models and achieve enhancements in robustness. Furthermore, our method achieves the best robust accuracy for CIFAR10, CIFAR100, and ImageNet datasets on the RobustBench leaderboard.
1 Introduction
Deep neural networks (DNNs) have demonstrated promising results across various tasks [1, 2], prompting concerns about AI security as these networks are increasingly deployed in our daily lives. A single erroneous prediction could lead to catastrophic consequences. For example, the Overload attack can significantly inflate the inference time of detecting objects for self-driving systems [3], while even minor typos in input prompts can cause large language models (LLMs) to produce unexpected responses [4].
The focus of this paper is to design robust DNNs that can defend against adversarial attacks, which aim to create perturbations in inputs that are imperceptible to humans but can mislead DNNs. Previous studies have revealed the existence of adversarial examples in diverse domains, such as image pixels [5], audio data [6], and textual content [7]. Consequently, exploring the vulnerabilities of DNNs and developing theoretically grounded explainable AI is crucial for ensuring the reliability of DNN-based applications.
Adversarial training [8] has proven to be an effective strategy for enhancing the robustness of DNNs. It achieves this by generating adversarial examples on the fly during the training phase and optimizing the model’s weights to minimize the losses caused by these examples. Recent studies have shown that robustness can be further improved by introducing additional examples from other datasets [9] or using generative models [10, 11] to cover low-frequency data. Despite the promising improvements in robustness, training costs increase significantly due to the demand for additional data, which can be up to 20 to 100 times larger than the original dataset. This poses a trade-off between training cost and robustness. The concern over high computational costs becomes a significant obstacle in deploying robust DNN-based applications, especially in fields like medicine, autonomous driving systems, and other areas where human lives are at stake.
In this paper, we explore how robustness is certified by the theorem of Lipschitz continuity, which theoretically gauges how much outputs are amplified by the perturbations. However, we argue that the set of observed data is finite and cannot cover the entire real data space, leading to an overestimation of the Lipschitz constant derived from the theorem. Therefore, we propose an algorithm that can remap the input domain into a constrained range, resulting in a Lipschitz constant for the modified function that is less than or equal to the Lipschitz constant of the original function, thus potentially enhancing robustness. Our key contributions are outlined as follows:
  • We introduce the concept of the empirical Lipschitz constant, which can more precisely reflect the robustness of the corresponding observed data. Compared with the original definition of Lipschitz constant, the empirical value is derived from a set of observed data, thereby eliminating the influence of space that is never drawn from real data. As illustrated in Figure 1, we prove that any function that can be formulated as a linear system, when combined with our proposed function to remap the input domain of a specific layer to a constrained range, can reduce its empirical Lipschitz constant, resulting in better robustness.
  • The proposed function can enhance the robustness of adversarially trained models with minimal additional costs. Specifically, it introduces only one parameter, the value of which can be determined by scanning an observed data once without the need for re-training or fine-tuning.
  • The experimental results suggest that our method can be combined with various existing methods and gain robustness improvements. Besides, our method achieves the best robust accuracy against adversarial examples generated by AutoAttack [12], a state-of-the-art ensemble attack, for CIFAR10, CIFAR100, and ImageNet datasets on the RobustBench leaderboard [13]. By assessing accuracy against adaptive attacks, transfer attacks, and evaluation methods for validating obfuscated gradients [14, 15], we believe that the proposed algorithm should not cause robustness to be overestimated.
Figure 1: The empirical Lipschitz constant of specific layers that can be represented by linear systems, such as convolutional or fully connected layers, can be reduced by remapping their input domain to a constrained range.
The rest of this paper is organized as follows. Section 2 introduces the background on adversarial attacks and adversarial training. Section 3 presents the theoretical proof of how robustness is enhanced by manipulating the domain of linear functions and introduces the proposed algorithm. Section 4 shows the experimental results, including the comparisons among related works, and ablation studies on various hyper-parameters, combination with different activation functions and gradient masking verification. The last section is our conclusion.
2 Related Works
2.1 Adversarial Attacks
Adversarial attacks aim to inject tiny perturbations into inputs, causing victim DNNs to output incorrect predictions with high confidence [16]. These attacks have been observed in numerous vision applications [17, 18, 19, 20]. Furthermore, these tiny perturbations can be embedded not only in image pixels but also in textual contexts [21, 22], audio space [23], and other fields [24]. Some research has shown how adversarial attacks threaten real applications [25, 26, 27, 28]. Investigating the vulnerability of DNNs and theoretically avoiding adversarial examples when optimizing model weights or designing architecture is an ongoing challenge.
Depending on the amount of information the attacker has access to, adversarial attacks can be divided into two types: white-box attacks and black-box attacks. For white-box attacks, all information about the victim models is public. Attackers can craft adversarial examples through the gradient direction, which is usually opposite to the direction in which the model weights were optimized during the training phase [8, 12]. Although this scenario is unrealistic, this type of research could lead to the design of more reliable models in the future. Conversely, for black-box attacks, the only information leaked to the attacker is the output prediction of the victim model. Adversarial examples can be generated by random search [29], discrepancies in outputs [30], or transferability from models with similar architectures [31, 32]. The purpose of black-box attacks is to study the risk of the victim models being attacked in real-world application scenarios.
2.2 Defensive Strategies
Adversarial training is a defensive strategy that aims to find optimal weights against adversarial attacks [8, 33, 34]. It achieves this by generating adversarial examples on the fly during the training phase and optimizing the model’s weights to minimize the losses caused by these examples. Despite the superior robustness achieved by adversarial training, the associated training costs of adversarially trained models are generally ten times more expensive than those of models trained utilizing a standard policy. The concern over high computational costs becomes a significant obstacle in deploying DNN-based applications.
Balancing between training cost and robustness is a challenge for adversarial training. Fast adversarial training has been proposed for applications pursuing higher robustness under a limited budget [35, 36]. However, numerous adversarial examples cannot be drawn from these approaches, potentially leading to catastrophic overfitting, where robust accuracy significantly decreases without warning signs [37]. On the contrary, some studies attempted to refine robustness by introducing additional examples from other datasets [9] or using generative models [10, 11]. Alternatively, another line of research has demonstrated that the removal of partial adversarial examples does not compromise robust accuracy, addressing the issue of unaffordable training costs [38, 39].
Despite the potential of adversarial training to enhance model robustness, budgetary constraints often limit the scope of their crafting to one or two specific attack types during the training stage. This restricted approach may inadvertently render adversarially trained models susceptible to novel, unseen attacks. As an alternative, Lipschitz-based certified training offers a theoretical framework for ensuring an upper bound on prediction errors [40, 41, 42]. However, it is important to acknowledge that these training methods often suffer from scalability issues.
3 Methodology
3.1 Motivation
In this paper, we approach robustness from a theoretical perspective, aiming to demonstrate that all risks posed by adversarial examples are limited while minimizing the additional costs associated with improving robustness. Our evaluation is conducted under the white-box scenario, where the target model is capable of defending against various types of known adversarial attacks, including white-box attacks [8, 18, 12], black-box attacks [43], and transfer attacks [44, 45]. Additionally, we conduct a set of experiments to verify that gradient masking [14] does not occur in our method and to ensure that robustness is not overestimated.
3.2 Lipschitz Continuity
To achieve our goal, we introduce a quantitative metric known as the Lipschitz constant, which gauges how much outputs are amplified by the perturbations within the input domain. The mathematical definition is as follows, a function $f:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ is globally Lipschitz continuous if there exists an constant $K\geq 0$ such that
$D_{f}(f(x_{1}),f(x_{2}))\leq KD_{x}(x_{1},x_{2})\quad\forall x_{1},x_{2}\in\mathbb{R}^{m},$ (1)
where $D_{x}$ is a metric on the domain of $f$; $D_{f}$ is a metric on the range of $f$; and $x_{1}\neq x_{2}$. For a DNN, it can be considered as a composite function:
$F(x)=(f_{1}\circ f_{2}\circ\dots\circ f_{L})(x),$ (2)
where $f_{i}$ is the function of $i$-th layer. If there exists a Lipschitz constant for each individual layer, we can derive an upper bound of the Lipschitz constant for the victim model as follows,
$K_{F}\leq\prod_{i=1}^{L}K_{i},$ (3)
where $K_{i}$ is the Lipschitz constant of $f_{i}$.
By defining adversarial examples $x^{\text{adv}}$ within a $\epsilon$-ball centered at an image $x$ as the inputs of (1), we can assess the impact caused by adversarial examples. Therefore, the Lipschitz constant serves as a bridge that connects the design of robust models with the measurement of risks posed by adversarial examples. A small Lipschitz constant for the victim model implies that the increase in loss is minimal, indicating a higher ability to resist adversarial attacks. Consequently, the objective of this paper is to lower the upper bound of Lipschitz constant for the given models.
As indicated by previous studies [46, 47], Lipschitz constant of the given model defined in (3) can be minimized by reducing the output discrepancy of individual linear layers. Under the $L_{2}$ norm, we have
$\frac{||f(x^{\text{adv}})-f(x)||_{2}}{||x^{\text{adv}}-x||_{2}}=\frac{(||Wx^{\text{adv}}+b)-(Wx+b)||_{2}}{||\delta||_{2}}=\frac{||W\delta||_{2}}{||\delta||_{2}},$ (4)
where $W$ is the weight matrix; and $\delta$ is the distance between $x^{\text{adv}}$ and $x$. Therefore, the original optimization problem of minimizing Lipschitz constant is transformed into the following minimization problem:
$\min_{W}\max_{\delta\neq 0,\delta\in\mathbb{R}^{m}}\frac{||W\delta||_{2}}{||\delta||_{2}}=\min_{W}\sigma_{\text{max}}(W),$ (5)
where $\sigma_{\text{max}}(W)$ represents the largest singular value of the matrix $W$. Notably, there is a relation to eigenvalues:
$\sigma^{2}_{i}(W)=\lambda_{i}(WW^{\dagger})=\lambda_{i}(W^{\dagger}W),$ (6)
where $W^{\dagger}$ is the conjugate transpose of $W$. Each singular value of the matrix $W$ is the square root of the eigenvalue of the matrices $WW^{\dagger}$ or $W^{\dagger}W$. In other words, minimizing $\lambda_{\text{max}}(WW^{\dagger})$, the largest eigenvalue of the matrices, can achieve the same objective.
Rather that minimizing the objective directly, Gershgorin circle theorem provides an alternative solution to estimate the robustness of the given linear system.
Theorem 1.
(Gershgorin Circle Theorem) For an $m\times m$ matrix $A$ with entries $a_{ij}$, each eigenvalue of $A$ is in at least one of the disk:
$R_{i}=\{z\in\mathbb{C}:|z-a_{ii}|\leq\sum_{i\neq j}|a_{ij}|\}\quad\mathrm{for}\quad i=\{1,2,\ldots,m\}.$ (7)
Theorem 1 indicates each row vector can be represented as a disk which is centered at the diagonal entry $a_{ii}$ and whose radius is the sum of the off-diagonal entries $a_{ij}$. For any layer which can be represented by a linear system, such as convolutional or fully connected layers, robustness can be improved by shrinking the radius of the disk with the largest eigenvalue.
3.3 Forged Function
(a) ConvNet
(b) Transformer
Figure 2: Insertion points of the forged function. In ConvNets, it is inserted into the residual blocks, while in Transformers, it is inserted into the MLP layers.
We argue that the largest singular value provides a loose bound for the Lipschitz constant. To precisely reflect the robustness of the corresponding observed data, we define the empirical Lipschitz constant that eliminates the influence of space that is never drawn from real data.
Definition 1.
Empirical Lipschitz constant:
$\max_{\delta\neq 0,x\in\mathcal{S}}\frac{||Wx||_{2}}{||x||_{2}}\quad\forall x\in\mathcal{S},$ (8)
where $\mathcal{S}$ is an observed dataset. As can be seen, the empirical Lipschitz constant on the finite dataset is less than or equal to its Lipschitz constant derived from the theorem.
Based on Definition 1, we can build robust models by manipulating the output ranges of individual layers, thereby restricting the input domain of the next layer. If input vectors do not align with the direction of the eigenvector with the largest eigenvalue, the empirical constant should be bounded. Therefore, we proposed a forged function defined as follows:
$f^{\text{forge}}(x)=\begin{cases}0\quad&\text{if}\quad|x|\leq c^{\text{th}}_{i},\\ x\quad&\text{otherwise},\end{cases}$ (9)
where $c^{\text{th}}_{i}$ is a threshold for the $i$-th layer. Compared with the original functions, the range of the forged function is suppressed if its value is less than the threshold. When $c^{\text{th}}_{i}$ is set to 00, the forged function degrades into the original function.
The forged function aim to reduce the empirical Lipschitz constant of the layers that can be represented as linear systems by remapping the input domain of these layers into a constrained set. Figure 2 provides a visual representation of potential insertion points for the forged function, while maintaining the integrity of other layers. For the ResNet architecture, the forged function is placed before the convolutional layers in each residual block. Similarly, for vision transformer architectures, the structure of MLP layers is adapted to seamlessly integrate the forged function.
Here is the proof that the largest eigenvalue can be shrunk by the forged function. Let $W$ be the weight of the target layer, which can be represented by an $m\times n$ matrix, and t be the input vector. Without loss of generality, we assume that $A=W^{\dagger}W$ and $f^{\text{forge}}(\textbf{t})$ is defined as:
$f^{\text{forge}}(t_{i})=\begin{cases}0\quad i\leq k\\ t_{i}\quad\text{otherwise},\end{cases}$ (10)
where $t_{i}$ is the $i$-th element of t and $k$ is a positive number.
Lemma 2.
There exists a matrix $A^{\prime}$ whose largest eigenvalue, $\lambda_{\text{max}}(A^{\prime})$, is less than or equal to the largest eigenvalue of $A$, $\lambda_{\text{max}}(A)$, if
$Af^{\text{forge}}(\textbf{t})=A^{\prime}\textbf{t}.$ (11)
Proof.
Since the first $k$ entries of the vector t are replaced with zeros, above condition can be achieved by replacing the corresponding column vectors of the matrix $A$ with zero vectors. Therefore, the entries of $A^{\prime}$ are formulated as
$a^{\prime}_{ij}=\begin{cases}0\quad j\leq k\\ a_{ij}\quad\text{otherwise}.\end{cases}$ (12)
The matrix $A$ is a positive semidefinite matrix, implying that the diagonal entries are non-negative. Moreover, with the entry representation of $A^{\prime}$ in (12), we observe that modifications are only applied to the first $k$ columns, while the rest remain unchanged. Combining the Gershgorin Circle Theorem, we know that the centers of the first $k$ disks of the matrix $A^{\prime}$ are shifted towards zero. Additionally, the radii of all disks, the absolute values of the off-diagonal entries in $A^{\prime}$, are shrunk. Consequently, the upper bound of the largest eigenvalue of the matrix $A^{\prime}$ is tighter compared to that of the original matrix $A$. ∎
Notably, the outputs of $f^{\text{forge}}(\textbf{t})$ vary depending on the inputs, resulting in each input having its own $A^{\prime}$. The upper bound of the largest eigenvalue of each matrix $A^{\prime}$ must be not greater than the largest eigenvalue of the matrix $A$. With Lamma 2, a precise upper bound of the largest eigenvalue can be obtained by feeding a set of observed images. On the contrary, there might be cases in which solving the minimization problem in (5) leads to the theoretical largest eigenvalue being minimized, but the empirical Lipschitz constant remains unchanged.
The choice of a proper $c^{\text{th}}_{i}$ is a crucial factor in reducing the largest eigenvalue. In this paper, we propose obtaining the value of $c^{\text{th}}_{i}$ through the following equation:
$c^{\text{th}}_{i}=c^{r}\max(F_{1\rightarrow i}(x))\quad\forall x\in\mathcal{S},$ (13)
where $\mathcal{S}$ can include all or a subset of images in the training set, $c^{r}$ is a positive number and $F_{1\rightarrow i}(x)$ represents the output of the $i$-th layer. Specifically, each layer has its own $c^{\text{th}}_{i}$, but they share the same hyper-parameter $c^{r}$. Algorithm 1 specifies the implementation details of the forged function. The variable $b$ is used to store the maximum value that appeared in $\mathcal{S}$, as defined in (13), and is initialized during construction. Similar to the implementation of the batchnorm layer, the behavior is depended on the mode configuration. When the mode is set to tracking mode, the variable $b$ is updated accordingly, and the input is set to the output without any modification. Conversely, when the mode is set to inference mode, the value of $b$ is frozen, but the input is updated as defined by (9). By default, the mode is set to inference, and the values of $b$ and $c^{r}$ are zero, respectively. As a result, the set $\mathcal{M}$ is empty, and the algorithm is degraded to the identical function.
It is worth emphasizing that by feeding all images in the set $\mathcal{S}$ once in track mode beforehand, the value of $c^{\text{th}}_{i}$ can be obtained. The elements satisfying the constraints are appropriately deactivated during inference. Notably, this operation does not necessitate gradient computations and incurs minimal time consumption, typically only a few minutes, even when executed on commonly used GPUs. In comparison to adversarial training, this process is nearly cost-free. The overall procedure shares many similarities with post-pruning techniques. Nevertheless, we posit that the proposed function is very similar to the ReLU function, as it suppresses the output values within a specific range, but the defined range in the forged function is adaptive to the observed dataset.
Algorithm 1 Forged Function
1:  require: Input x, Mode $m$
2:  if $m$ is tracking mode then
3:     $b=\max(b,\textbf{x})$
4:  else
5:     =$\mathcal{M}=\left\{x\middle|\text{abs}(x)\leq c^{r}b\right\}$
6:     for all $s\in\mathcal{M}$ do
7:        $s=0$
8:     end for
9:  end if
10:  return x
4 Experiments
4.1 Setup
We evaluated the performance on CIFAR10, CIFAR100, and ImageNet datasets under the white-box scenario with an $L_{\infty}$ norm. To ensure comparability of results, we assessed robustness using AutoAttack [12], For CIFAR10 and CIFAR100 datasets, $\epsilon$ is set to $8/255$, while for the ImageNet dataset, $\epsilon$ is set to $4/255$. The model weights are publicly accessible from RobustBench. The ablation study involves exploring the selection of the optimal $c^{r}$, the combination of various models trained from different techniques, the verification of gradient masking, and assessments of certified adversarial robustness via randomized smoothing. Due to the page limit, the full experimental results of the ablation study are listed in Appendix.
4.2 White-box Evaluation
4.2.1 Performance Analysis on CIFAR10 and CIFAR100 Datasets
Table 1: The results of top-3 competitors on Robustbench.
(a) CIFAR10 dataset
# Method accnat accAA
* [11] + Ours 93.20 71.70
1 [48] 93.27 71.07
2 [11] 93.25 70.69
3 [49] 95.19 69.71
(b) CIFAR100 dataset
# Method accnat accAA
* [11] + Ours 74.97 44.00
1 [11] 75.22 42.67
2 [49] 83.08 41.80
3 [50] 73.85 39.18
The model used in this study is based on WRN-70-16 architecture with SiLU function while generative data were involved during the training phase. The value of $c^{\text{th}}_{i}$ is obtained by feeding all images from the training set without any augmentation, and $c^{r}$ was set to $2^{-8}$ for this experiment. Tables 1(a) and 1(b) summarize the top-3 competitors on Robustbench for CIFAR10 and CIFAR100 datasets, respectively, where ##\## represent the rankings, our results are marked by the asterisk (*), accnat and accAA denote the accuracy against clean data and adversairal examples generated by AutoAttack, respectively.
As can be seen, our method combined with WRN-70-16 with SiLU function gains improvement in robustness by at least 0.9% and achieves the best results on Robustbench for both datasets. Nevertheless, standard accuracy (accnat) is decreased. Many factors might affect the results. For example, the single additional hyper-parameter introduced in this study might not provide sufficient granularity to fit all layers in the target model.
4.2.2 Performance Analysis on ImageNet Dataset
Table 2: The results of top-3 competitors for ImageNet dataset on Robustbench.
# Method Architecutre accnat accAA
* [51] + Ours Swin-L 78.88 60.04
1 [51] Swin-L 78.92 59.56
2 [49] ConvNeXtV2-L + Swin-L 81.48 58.50
3 [51] ConvNeXt-L 78.02 58.48
In this experiment, we utilized the Swin [52] model architecture, a variant of transformers. However, scanning the approximately 1.2 million training images provided by the ImageNet dataset to determine the value of the hyper-parameter introduced in the forged function defined in (13) might take a long time. Alternatively, we randomly selected about 5,000 images as the observed images to determine the value of the hyper-parameter. Ideally, determining the optimal choice of $c^{r}$ requires conducting an ablation study to explore the relationship between the chosen $c^{r}$ and robust accuracy on a validation set. To accelerate this procedure, we first seek a value of $c^{r}$ with the highest standard accuracy. The candidate values are selected in a small range centered around this value.
Tables 2 lists the top-3 competitors on Robustbench for ImageNet dataset, including ranking, architecture, standard accuracy, and robust accuracy against AutoAttack. The experimental results demonstrate that the Swin-L model with GELU combined with our method can obtain improvements in robust accuracy and achieve the best result while standard accuracy has a tiny drop. This finding verifies that our method can be applied to both convolutional and fully connected layers.
4.2.3 Combination with Various Models
The experiment aims to assess the generability of the proposed function on adversarial trained models with identical architecture but from various training strategies and to evaluate the potential cost reduction of adversarial training. We integrated the proposed approach with partial models selected from RobustBench, whose weights are obtained directly from the official without any modifications, and also included a model trained by TRADES [53] as a baseline for CIFAR10 dataset. The selected models were trained using different techniques, such as adding perturbations in internal layers, retrieving information using knowledge distillation, reducing inefficient training data, or involving additional images from generated models or another dataset. Except for the model used in RST-WAP, which is WRN-28-10, the model architecture we utilized is WRN-34-10 with ReLU, as it is the most popular network on the RobustBench leaderboard [13].
For the white-box evaluation, the value of $c^{r}$ is set to $2^{-7}$. Table 3(a) and 3(b) present standard and robust accuracy of models integrated with our method for CIFAR10 and CIFAR100 dataset, respectively. In these tables, the column Original indicates the original results reported by RobustBench, and the column Original+Ours demonstrates the results of the proposed method. As indicated in these tables, for CIFAR10 dataset, the proposed method enhances robust accuracy by more than $2\%$ for RST-AWP, DefEAT, and LTD models, while other models receive approximately $1$ to $1.5\%$ improvement in robustness. Similarly, for CIFAR100 dataset, these models meet at least a $1\%$ increase in robustness. The empirical results prove that the resilience of existing models against adversarial attacks can be improved by Lemma 2. We believe that the proposed solution is general as it achieves great success in models incorporating different training techniques.
Another advantage of the proposed method that we would like to highlight is that the cost of our approach can almost be ignored compared to the cost of adversarial training as the cost involves only a single pass scan of a set of images to determine the hyper-parameter $c^{th}$. This implies that these models can enhance robustness for free. Specifically, LefEAT can achieve a robust accuracy of 57.30% by removing inefficient training data. By combining LefEAT model with our approach, a robust accuracy of 59.55% can be achieved, which is comparable to RST-AWP (60.04%). However, RST-AWP introduces more images from another dataset, resulting in a higher cost in each epoch. Similarly, for the CIFAR100 dataset, DefEAT with our proposed method achieves a robust accuracy of 32.11%, which is better than EffAug (31.85%), which involves more complex data augmentation during the training stage. This aligns with the suggestion by DefEAT that some data can be removed without hurting robustness. Holistically, we believe that our approach might provide a hint during the late phase of adversarial training to drop inefficient weights, resulting in further cost savings or enhanced resilience.
Another interesting observation presented in these tables is that standard accuracy is also improved for all models on both CIFAR10 and CIFAR100 datasets. Although this phenomenon cannot be explained by Lemma 2, we believe that the output range between ReLU and the proposed functions has a high similarity. Consequently, this arrangement can maintain the accuracy of clean data.
Table 3: Standard and robust accuracy of models integrated with our method.
(a) CIFAR10 dataset
Method Original Original+Ours accnat accAA accnat accAA RST-AWP [54] 88.25 60.04 89.50 62.76 DefEAT [39] 86.54 57.30 87.40 59.55 LTD [55] 85.21 56.94 85.98 59.25 AWP [54] 85.36 56.17 86.19 57.85 TRADES[53] 85.34 52.86 85.78 53.80
(b) CIFAR100 dataset
Method Original Original+Ours accnat accAA accnat accAA EffAug [56] 68.75 31.85 69.14 32.57 DKLD [50] 64.08 31.65 64.26 32.58 DefEAT [39] 64.32 31.13 66.42 32.11 LTD [55] 64.07 30.59 64.29 31.95 AWP [54] 60.38 28.86 60.63 29.72
4.2.4 Gradient Masking Verification
Previous studies suggest that the resilience of models might be unintentionally overestimated [14, 15]. The proposed function in (9) suppresses values to zero if the condition is satisfied. One might argue that this property could unintentionally cause obfuscated gradients, resulting in gradient attacks being unable to efficiently produce adversarial examples. Therefore, to verify that the proposed method does not encounter the gradient masking issue, we should conduct more experiments from the following aspects:
  1. 1.
    White-box attacks should be better than black-box attacks.
  2. 2.
    Iterative attacks should have better performance than one-step attacks.
  3. 3.
    Robust accuracy should gradually decrease to zero when the radius of $\epsilon$-ball increase.
  4. 4.
    The modified model should defense against adversarial examples generated by the original models.
  5. 5.
    Certified robustness that conducted by random smoothing [57].
The first item has been examined by AutoAttack as it involves three white-box attacks and one black-box attack. By compared robust accuracy shown in Table 1(a), 1(b) and 2, the models combined with the proposed method perform better robust accuracy than the original models. It indicates black-box cannot produce more adversarial examples.
The full experimental results for the rest of experiments can be found in the appendix. The results demonstrate that the proposed algorithm does not violate any of the above rules and the certified robustness improves by our method across most settings. From the evidence, we believe that the proposed method does not encounter the gradient masking problem among different hyper-parameters and various models on CIFAR10 and CIFAR100 datasets.
5 Conclusion
In this paper, we recap how robustness is certified by the theorem of Lipschitz continuity. We introduce the concept of the empirical Lipschitz constant, which minimizes the influence of the space not drawn from real data, resulting in a precise estimation of the robustness of the corresponding observed data. We prove that by remapping the input domain of a specific layer to a constrained range, the Lipschitz constant can be shrunk, leading to better robustness. The proposed function introduces only one parameter, the value of which can be determined by scanning the training data once, without re-training or fine-tuning. Compared with adversarial training, the proposed method is almost cost-free. The experimental results suggest that our method can be combined with various existing methods and achieve robustness improvements, and no gradient masking occurs in our algorithm. Furthermore, our method can achieve the best robust accuracy for CIFAR10, CIFAR100, and ImageNet datasets on the RobustBench leaderboard.
Numerous future directions merit exploration. Firstly, due to the property of maximization, the proposed function might easily be influenced by outliers. Designing a better function is an interesting research topic. Secondly, exploring the combination with various activation functions, different model architectures or large-scale datasets would be beneficial. Lastly, it is worth investigating to understand the theoretical reasons why our proposed function improves standard accuracy.
Appendix A Ablation Study
A.1 Hyper-parameter Selection
This experiment investigates how the choice of hyper-parameter $c^{r}$ influences standard accuracy and robust accuracy. Since most models are represented in 16 bit format, and the widths of fraction bit for FP16 format defined by IEEE-754 standard and BFloat are 10 and 7 bits, respectively, truncated errors might easily occur when performing addition on two numbers with a magnitude difference of $2^{8}$ or higher. On the other hand, when $c^{r}$ is set to $2^{-5}$, all models experience a significant drop in standard accuracy, and there is meaningless in evaluating robustness at this configuration. We suggest that the candidates of $c^{r}$ are $2^{-8}$, $2^{-7}$ and $2^{-6}$.
The results on CIFAR10 and CIFAR100 are presented in Table 5 and Table 5, respectively. Moreover, the results of accuracy against CW attack on $L_{\infty}$ norm for CIFAR10 and CIFAR100 datasets are presented in Tables 6(a) and 6(b), respectively. As can be seen, when $c^{r}$ is set to $2^{-8}$, all models achieve better standard accuracy and robust accuracy. Additionally, the results for all models with $c^{r}=2^{-7}$ are surpassed by those when $c^{r}$ is set to $2^{-8}$. Robust accuracy can be further enhanced by setting $2^{-6}$, while standard accuracy might drop compared to the original. The results suggest that $c^{r}=2^{-7}$ is a solution that balances standard accuracy and robustness. Nevertheless, when robustness is a major concern, $c^{r}=2^{-6}$ is a better choice.
Intuitively, we expect that standard accuracy gradually decreases when the value of $c^{r}$ increases. The phenomenon can be observed when $c^{r}$ is $2^{-6}$ or higher but two counterexamples are reported in the ablation study when setting $c^{r}$ to $2^{-7}$ and $2^{-8}$. A possible explanation is that the optimizer becomes stuck in a saddle area, as ReLU is non-differentiable at the zero point. This might cause the gradient direction to become stuck in an oscillation when values are close to zero. By shifting those values to zero, antagonistic effects among different feature maps, filters, or channels are accidentally mitigated. However, further investigation and evidence are needed to support this conjecture.
We argue that any function that satisfies the conditions defined in (11) can shrink the largest eigenvalue. There might be another function that can perform better than the proposed one. Besides, the hyper-parameter is determined by choosing the maximum value appearing in the dataset.
Table 4: Ablation study of selecting optimal $c^{r}$ for CIFAR10 dataset.
Method RobustBench $c^{r}=2^{-8}$ $c^{r}=2^{-7}$ $c^{r}=2^{-6}$
accnat accAA accnat accAA accnat accAA accnat accAA
RST-AWP 88.25 60.04 88.82 60.96 89.50 62.76 87.88 61.96
DefEAT 86.54 57.30 86.88 57.81 87.40 59.55 84.59 61.08
LTD 85.21 56.94 85.28 57.28 85.98 59.25 85.59 60.63
AWP 85.36 56.17 85.80 56.53 86.19 57.85 84.55 59.21
TRADES 85.34 52.86 85.57 52.97 85.78 53.80 85.49 55.37
Table 5: Ablation study of selecting optimal $c^{r}$ for CIFAR100 dataset.
Method RobustBench $c^{r}=2^{-8}$ $c^{r}=2^{-7}$ $c^{r}=2^{-6}$
accnat accAA accnat accAA accnat accAA accnat accAA
EffAug 68.75 31.85 68.81 32.00 69.14 32.57 68.44 33.64
DKLD 64.08 31.65 64.10 31.77 64.26 32.58 63.50 33.87
DefEAT 65.89 30.57 66.12 31.11 66.42 32.46 65.06 34.07
LTD 64.07 30.59 64.29 31.13 64.29 31.95 64.18 34.04
AWP 60.38 28.86 60.18 29.10 60.63 29.72 60.71 30.82
Appendix B Full Experimental Results of Gradient Masking Verification
Table 6: The robust accuracy against CW attack on $L_{\infty}$ norm.
(a) CIFAR10 dataset
Method Origin $c^{r}$
$2^{-8}$ $2^{-7}$ $2^{-6}$
RST-AWP 58.98 61.84 68.24 80.92
DefEAT 56.92 58.02 61.06 65.56
LTD 58.12 58.56 60.50 64.86
AWP 56.84 57.34 60.58 66.50
TRADES 56.10 56.52 58.18 63.62
(b) CIFAR100 dataset
Method Origin $c^{r}$
$2^{-8}$ $2^{-7}$ $2^{-6}$
EffAug 37.40 37.70 38.70 43.00
DKLD 37.50 38.06 39.38 44.20
DefEAT 36.90 37.56 39.82 44.30
LTD 36.66 37.32 38.86 43.44
AWP 34.56 35.20 35.94 40.40
Table 7(a) and 7(b) present the robust accuracy against adversarial examples generated by the original models on CIFAR10 and CIFAR100 datasets, respectively. As observed, none of the models showed lower robust accuracy than the original model. It indicates that adversarial examples can be efficiently crafted by utilizing the gradients from the victim models.
Table 8 and 9 presents the robust accuracy against FGSM and PGD attacks among different radii of the $\epsilon$-ball on the CIFAR10 and CIFAR100 datasets, respectively. As observed, the robust accuracy against FGSM, a one-step attack, is always higher than the robust accuracy against PGD, an iterative attack. This implies that the gradient is reliable, allowing the PGD attack to adjust the gradient direction multiple times to find adversarial examples. Additionally, we observe that the robust accuracy against PGD attacks for all models gradually decreases to zero as the radius of the $\epsilon$-ball increases. This indicates that the quality of gradients is preserved, enabling PGD attacks to move the gradient toward examples not in the observed distribution.
Figure 3 illustrates the certified robustness achieved by random smoothing for various models on the CIFAR10 dataset, where Original refers to the certified robustness of the original model, while Ours denotes the robustness of the model combined with the proposed method. As can be seen, our method brings slight improvements in robustness, except for the AWP model. These results demonstrate that our algorithm does not suffer from the gradient masking issue. However, the empirical Lipschitz constant is derived from the observed data. As the input distribution drawn from random smoothing and the observed data might have discrepancies, this could result in fluctuations in robustness.
(a) RST-AWP
(b) DefEAT
(c) LTD
(d) AWP
Figure 3: Certified robustness that conducted by random smoothing.
Table 7: The robust accuracy against adversarial examples generated by the original models.
(a) CIFAR10 dataset
Method Origin $c^{r}$
$2^{-8}$ $2^{-7}$ $2^{-6}$
RST-AWP 60.04 62.10 65.10 70.53
DefEAT 57.30 58.37 60.39 66.10
LTD 56.94 58.71 61.63 66.47
AWP 56.17 57.49 59.74 65.58
TRADES 52.86 55.55 55.09 58.68
(b) CIFAR100 dataset
Method Origin $c^{r}$
$2^{-8}$ $2^{-7}$ $2^{-6}$
EffAug 31.85 32.87 35.08 40.04
DKLD 31.65 32.91 35.04 40.58
DefEAT 30.57 31.82 33.94 40.67
LTD 30.59 32.05 34.07 39.11
AWP 28.86 29.88 32.18 36.67
Table 8: The robust accuracy against FGSM and PGD attacks among different radii of $\epsilon$-ball on CIFAR10 dataset.
Method $c^{r}$ Attack $\epsilon$
$\frac{1}{255}$ $\frac{2}{255}$ $\frac{4}{255}$ $\frac{8}{255}$ $\frac{16}{255}$ $\frac{32}{255}$ $\frac{64}{255}$ $\frac{96}{255}$
RST-AWP $2^{-8}$ FGSM 88.28 86.94 83.80 75.12 57.23 34.04 18.80 19.03
PGD 86.78 84.47 79.03 66.03 34.02 2.01 0.01 0.0
$2^{-7}$ FGSM 89.46 88.28 85.64 77.62 60.60 35.91 19.39 20.68
PGD 88.03 85.88 80.89 69.24 38.27 3.08 0.1 0.0
$2^{-6}$ FGSM 87.70 86.63 84.38 77.91 60.93 33.48 16.12 18.81
PGD 86.38 84.69 81.19 73.72 52.24 11.89 0.19 0.0
DefEAT $2^{-8}$ FGSM 86.38 85.40 81.98 72.73 53.28 30.34 18.13 19.65
PGD 84.52 82.07 76.51 63.71 33.87 1.76 0.0 0.0
$2^{-7}$ FGSM 86.69 85.70 83.05 74.35 56.36 32.04 18.88 21.40
PGD 85.11 82.87 78.00 66.54 38.70 3.12 0.0 0.0
$2^{-6}$ FGSM 84.14 83.57 81.03 74.63 57.67 28.11 13.17 19.38
PGD 82.96 81.35 77.37 69.52 50.70 10.05 0.2 0.0
LTD $2^{-8}$ FGSM 84.94 83.87 81.15 72.80 55.45 33.06 18.44 17.48
PGD 83.13 80.68 75.53 63.52 34.81 2.64 0.0 0.0
$2^{-7}$ FGSM 85.48 84.67 82.24 74.10 57.41 35.53 17.57 17.50
PGD 83.88 81.88 77.00 65.38 28.57 3.95 0.0 0.0
$2^{-6}$ FGSM 85.06 84.28 82.21 75.77 60.79 34.39 14.34 15.42
PGD 83.91 82.00 78.33 69.82 49.95 11.63 0.21 0.0
AWP $2^{-8}$ FGSM 85.11 83.90 80.68 71.28 53.78 33.21 20.86 19.61
PGD 83.34 80.34 75.08 61.53 30.50 1.89 0.03 0.0
$2^{-7}$ FGSM 85.50 84.68 81.75 73.56 57.19 35.50 20.63 20.16
PGD 83.94 81.62 76.34 65.57 34.16 2.89 0.02 0.0
$2^{-6}$ FGSM 83.87 83.19 81.08 74.97 61.49 35.13 16.12 18.69
PGD 83.00 81.42 78.13 71.50 54.95 14.78 0.41 0.0
TRADES $2^{-8}$ FGSM 84.74 83.51 70.58 70.50 54.23 36.53 23.97 23.46
PGD 82.62 79.72 72.81 57.30 24.21 1.21 0.01 0.0
$2^{-7}$ FGSM 85.00 84.02 80.57 71.61 55.94 25.67 22.31 22.51
PGD 82.96 80.31 73.75 58.91 26.22 1.31 0.02 0.0
$2^{-6}$ FGSM 85.05 84.19 81.40 75.07 60.24 36.99 21.81 21.21
PGD 83.23 81.33 75.81 64.56 36.76 3.37 0.02 0.0
Table 9: The robust accuracy against FGSM and PGD attacks among different radii of $\epsilon$-ball on CIFAR100 dataset.
Method $c^{r}$ Attack $\epsilon$
$\frac{1}{255}$ $\frac{2}{255}$ $\frac{4}{255}$ $\frac{8}{255}$ $\frac{16}{255}$ $\frac{32}{255}$ $\frac{64}{255}$ $\frac{96}{255}$
EffAug $2^{-8}$ FGSM 68.02 65.96 60.36 49.65 33.90 17.39 7.11 5.77
PGD 64.92 61.04 52.82 39.37 17.53 1.81 0.0 0.0
$2^{-7}$ FGSM 68.41 66.56 61.84 51.83 36.59 18.34 7.02 6.70
PGD 65.66 62.02 54.58 41.55 19.70 2.26 0.0 0.0
$2^{-6}$ FGSM 67.84 67.25 64.65 57.97 44.23 22.20 8.26 8.85
PGD 66.38 64.20 59.63 50.89 33.98 7.27 0.17 0.0
DKLD $2^{-8}$ FGSM 63.47 61.94 58.15 48.86 34.39 17.73 6.26 3.66
PGD 60.71 57.14 50.44 38.14 17.38 1.97 0.0 0.0
$2^{-7}$ FGSM 63.55 62.31 58.65 50.37 36.49 18.69 6.36 4.29
PGD 61.06 57.84 51.51 39.99 19.71 2.37 0.0 0.0
$2^{-6}$ FGSM 63.26 62.77 60.41 55.18 43.86 21.17 5.50 4.97
PGD 61.67 59.62 55.83 48.06 33.51 7.55 0.17 0.0
DefEAT $2^{-8}$ FGSM 65.57 64.36 59.88 49.38 32.48 15.38 5.50 3.30
PGD 62.39 58.96 51.85 38.59 17.18 1.45 0.0 0.0
$2^{-7}$ FGSM 65.97 64.96 60.67 51.44 34.84 16.36 5.33 3.96
PGD 62.94 59.83 52.98 41.05 20.07 2.02 0.0 0.0
$2^{-6}$ FGSM 64.69 63.58 61.31 54.47 40.06 16.91 4.64 4.46
PGD 62.85 60.50 56.29 47.54 30.76 5.84 0.09 0.0
LTD $2^{-8}$ FGSM 63.59 62.35 58.10 48.86 33.11 16.78 5.92 3.20
PGD 61.06 57.65 50.70 38.21 18.21 1.98 0.0 0.0
$2^{-7}$ FGSM 64.05 62.87 59.02 50.16 34.90 17.27 5.54 3.11
PGD 61.51 58.32 51.84 39.89 20.32 2.26 0.0 0.0
$2^{-6}$ FGSM 63.62 62.98 60.96 54.96 41.51 19.78 4.69 3.12
PGD 61.90 59.68 55.23 46.49 29.13 5.74 0.05 0.0
AWP $2^{-8}$ FGSM 59.77 58.06 54.21 45.54 30.98 16.69 6.49 3.97
PGD 56.72 52.92 46.53 34.90 16.03 2.11 0.0 0.0
$2^{-7}$ FGSM 60.00 58.52 54.93 46.65 32.66 17.42 6.04 3.60
PGD 57.19 53.75 47.46 36.22 17.48 2.58 0.0 0.0
$2^{-6}$ FGSM 60.20 59.71 57.47 52.05 39.78 21.10 5.70 3.90
PGD 58.19 55.66 50.91 42.37 25.53 5.68 0.09 0.0
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claim is stated in the abstract and highlighted in the introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We leave some topics unexplored for further work as described in the conclusion.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The proof of Lemma 2 is shown in Section 2.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The experiment were conducted by the public leaderboard RobusBench. The process is standardized and hardware-independent. The results can be fare compared to the competitiers.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [N/A]
  24. Justification: The model architecture and pre-trained weights proposed in this work will be submitted to Robusbench. This information will be published once the results are certified by Robusbench.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The assessments utilized by RobustBench are standardized.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: The assessments utilized by RobustBench are standardized. The impact of Statistical variances have been included in the assessments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [N/A]
  39. Justification: The assessments utilized by RobustBench are hardware-independent. Experiments can be performed on any GPU, but powerful GPUs can complete the evaluation in shorter execution times.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This work follows the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: The main goal of this work is to design reliable models. This work is unlikely to have a negative social impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This models used in this works are existing public models. Therefore, this work poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The related works and RobustBench are correctly cited in this work.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This work does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This work does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This work does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@
 Abstract
-The security and robustness of deep neural networks (DNNs) have become increasingly concerning.
-This paper aims to provide both a theoretical foundation and a practical solution to ensure the reliability of DNNs.
-We explore the concept of Lipschitz continuity to certify the robustness of DNNs against adversarial attacks, which aim to mislead the network with adding imperceptible perturbations into inputs.
-We propose a novel algorithm that remaps the input domain into a constrained range, reducing the Lipschitz constant and potentially enhancing robustness.
-Unlike existing adversarially trained models, where robustness is enhanced by introducing additional examples from other datasets or generative models, our method is almost cost-free as it can be integrated with existing models without requiring re-training.
-Experimental results demonstrate the generalizability of our method, as it can be combined with various models and achieve enhancements in robustness.
-Furthermore, our method achieves the best robust accuracy for CIFAR10, CIFAR100, and ImageNet datasets on the RobustBench leaderboard.
+As deep neural networks (DNNs) are increasingly deployed in sensitive applications, ensuring their security and robustness has become critical.
+A major threat to DNNs arises from adversarial attacks, where small input perturbations can lead to incorrect predictions.
+Recent advances in adversarial training improve robustness by incorporating additional examples from external datasets or generative models.
+However, these methods often incur high computational costs, limiting their practicality and hindering real-world deployment.
+In this paper, we propose a cost-efficient alternative based on Lipschitz continuity that achieves robustness comparable to models trained with extensive supplementary data.
+Unlike conventional adversarial training, our method requires only a single pass over the dataset without gradient estimation, making it highly efficient.
+Furthermore, our method can integrate seamlessly with existing adversarial training frameworks and enhances the robustness of models without requiring extra generative data.
+Experimental results show that our approach not only reduces computational overhead but also maintains or improves the defensive capabilities of robust neural networks.
+This work opens a promising direction for developing practical, scalable defenses against adversarial attacks.
 1 Introduction
-Deep neural networks (DNNs) have demonstrated promising results across various tasks [1, 2], prompting concerns about AI security as these networks are increasingly deployed in our daily lives.
-A single erroneous prediction could lead to catastrophic consequences.
-For example, the Overload attack can significantly inflate the inference time of detecting objects for self-driving systems [3], while even minor typos in input prompts can cause large language models (LLMs) to produce unexpected responses [4].
-The focus of this paper is to design robust DNNs that can defend against adversarial attacks, which aim to create perturbations in inputs that are imperceptible to humans but can mislead DNNs.
-Previous studies have revealed the existence of adversarial examples in diverse domains, such as image pixels [5], audio data [6], and textual content [7].
-Consequently, exploring the vulnerabilities of DNNs and developing theoretically grounded explainable AI is crucial for ensuring the reliability of DNN-based applications.
-Adversarial training [8] has proven to be an effective strategy for enhancing the robustness of DNNs.
-It achieves this by generating adversarial examples on the fly during the training phase and optimizing the model’s weights to minimize the losses caused by these examples.
+Deep neural networks (DNNs) have achieved remarkable success across a wide range of tasks [29, 38], driving their adoption in increasingly sensitive applications.
+As a result, concerns over the security and robustness of these models have become more urgent, as even a single erroneous prediction can have serious consequences.
+For example, the Overload attack can significantly inflate the inference time of detecting objects for edge devices [12], and even minor typos in prompts can cause large language models to behave unpredictably [55].
+Among the most critical threats are adversarial attacks, which add imperceptible perturbations to inputs, leading DNNs to make incorrect predictions.
+These vulnerabilities have been demonstrated across various domains [40, 6, 31], highlighting the urgent need for more robust and explainable AI models.
+Adversarial training [34] has proven to be an effective strategy for enhancing the robustness of DNNs by generating adversarial examples during training and optimizing model weights to minimize the losses caused by these examples.
+Despite promising results, no certified models have been deployed commercially due to prohibitive computational requirements.
 Recent studies have shown that robustness can be further improved by introducing additional examples from other datasets [8] or using generative models [24, 43] to cover low-frequency data.
-Despite the promising improvements in robustness, training costs increase significantly due to the demand for additional data, which can be up to 20 to 100 times larger than the original dataset.
-This poses a trade-off between training cost and robustness.
-The concern over high computational costs becomes a significant obstacle in deploying robust DNN-based applications, especially in fields like medicine, autonomous driving systems, and other areas where human lives are at stake.
-In this paper, we explore how robustness is certified by the theorem of Lipschitz continuity, which theoretically gauges how much outputs are amplified by the perturbations.
-However, we argue that the set of observed data is finite and cannot cover the entire real data space, leading to an overestimation of the Lipschitz constant derived from the theorem.
-Therefore, we propose an algorithm that can remap the input domain into a constrained range, resulting in a Lipschitz constant for the modified function that is less than or equal to the Lipschitz constant of the original function, thus potentially enhancing robustness.
+However, these approaches require datasets 100 or more times larger than the original, which significantly escalates computational demands by orders of magnitude.
+This creates a critical trade-off between training cost and robustness, presenting a significant obstacle to deploying robust DNN-based applications commercially are at stake.
+The focus of this paper is to explore an alternative approach that reduces training cost while achieving robustness comparable to state-of-the-art adversarial training frameworks.
+To address this issue, we first revisit the theorem of Lipschitz continuity, which gauges how much outputs are amplified by perturbations.
+Unlike certified training [35] or Lipschitz-constrained methods [56] approaches that primarily aim to tighten upper bounds but often face scalability issues with larger models and datasets, our approach takes a fundamentally different direction.
+We prove that remapping the input domain of any layer that can be represented as a linear system to a constrained range results in a bounded Lipschitz constant and robustness improvements when specific conditions are met.
+The key insight is that the forged function that reshapes the input domain of a linear system should maintain high similarity between the output of the transformed layers and the original, according to specific metrics.
+We propose a data-driven algorithm that constructs the optimal function with only a single scan over the dataset and no need for gradient estimation, making it highly efficient.
+This approach stands in contrast to prior methods that typically reshape weight distributions through regularization or by introducing complex objectives, which increases the computational cost of seeking suitable hyper-parameters.
 Our key contributions are outlined as follows:
-  • We introduce the concept of the empirical Lipschitz constant, which can more precisely reflect the robustness of the corresponding observed data.
-Compared with the original definition of Lipschitz constant, the empirical value is derived from a set of observed data, thereby eliminating the influence of space that is never drawn from real data.
-As illustrated in Figure 1, we prove that any function that can be formulated as a linear system, when combined with our proposed function to remap the input domain of a specific layer to a constrained range, can reduce its empirical Lipschitz constant, resulting in better robustness.
-  • The proposed function can enhance the robustness of adversarially trained models with minimal additional costs.
-Specifically, it introduces only one parameter, the value of which can be determined by scanning an observed data once without the need for re-training or fine-tuning.
-  • The experimental results suggest that our method can be combined with various existing methods and gain robustness improvements.
-Besides, our method achieves the best robust accuracy against adversarial examples generated by AutoAttack [12], a state-of-the-art ensemble attack, for CIFAR10, CIFAR100, and ImageNet datasets on the RobustBench leaderboard [13].
-By assessing accuracy against adaptive attacks, transfer attacks, and evaluation methods for validating obfuscated gradients [14, 15], we believe that the proposed algorithm should not cause robustness to be overestimated.
-Figure 1: The empirical Lipschitz constant of specific layers that can be represented by linear systems, such as convolutional or fully connected layers, can be reduced by remapping their input domain to a constrained range.
+  • Cost-Effective Robustness Enhancement: We propose a method that enhance the robustness of existing adversarially trained methods with minimal additional costs.
+Compared to previous works, where improved robustness relied on introducing more generative data during the adversarial training phase, our method can be combined with existing models without re-training or fine-tuning.
+From a different perspective, for models with similar robustness, those integrated with our approach require significantly lower training cost.
+  • Theoretical Foundation: We provide a theoretical proof that, under specific conditions, remapping the input domain of any linear system to a constrained range leads to a bounded Lipschitz constant and improved robustness.
+This theoretical foundation distinguishes our work from previous Lipschitz-based approaches that focus primarily on constraining the network during training.
+  • Inference-Time Optimization: The proposed method is well-suited for inference-time model optimization and is almost cost-free.
+The function introduces only one parameter, the value of which can be determined by scanning observed data once without requiring gradient computations or fine-tuning.
+This makes our approach particularly valuable for practical applications where computational resources are limited.
+  • State-of-the-Art Results: Experimental results demonstrate that our method can be combined with various existing methods to achieve meaningful robustness improvements.
+Notably, models such as LTD [10] and DefEAT [11], when combined with our method, achieve performance on par with RST-AWP [45], a method that relies on extra data during training, against against adversarial examples generated by AutoAttack [17], a state-of-the-art ensemble attack, for CIFAR10, CIFAR100, and ImageNet datasets on the RobustBench leaderboard [18].
+This finding supports the hypothesis that certain data samples may be safely omitted without sacrificing robustness, reinforcing the potential of developing data selection algorithms for adversarial training to reduce training cost without compromising effectiveness.
 The rest of this paper is organized as follows.
@@
 Section 3 presents the theoretical proof of how robustness is enhanced by manipulating the domain of linear functions and introduces the proposed algorithm.
-Section 4 shows the experimental results, including the comparisons among related works, and ablation studies on various hyper-parameters, combination with different activation functions and gradient masking verification.
+Section 4 shows the experimental results, including the comparisons among related works.
+Due to page limitations, ablation studies on various hyper-parameters, combination with different activation functions and gradient masking verification are left in Appendix.
 The last section is our conclusion.
@@
 Some research has shown how adversarial attacks threaten real applications [47, 28, 20, 44].
-Investigating the vulnerability of DNNs and theoretically avoiding adversarial examples when optimizing model weights or designing architecture is an ongoing challenge.
-Depending on the amount of information the attacker has access to, adversarial attacks can be divided into two types: white-box attacks and black-box attacks.
-For white-box attacks, all information about the victim models is public.
-Attackers can craft adversarial examples through the gradient direction, which is usually opposite to the direction in which the model weights were optimized during the training phase [8, 12].
-Although this scenario is unrealistic, this type of research could lead to the design of more reliable models in the future.
-Conversely, for black-box attacks, the only information leaked to the attacker is the output prediction of the victim model.
-Adversarial examples can be generated by random search [29], discrepancies in outputs [30], or transferability from models with similar architectures [31, 32].
-The purpose of black-box attacks is to study the risk of the victim models being attacked in real-world application scenarios.
+Investigating the vulnerability of DNNs theoretically and designing robust DNNs is an ongoing challenge.
+Adversarial attacks can be classified into two types based on the amount of information the attacker has access to: white-box attacks and black-box attacks.
+In the case of white-box attacks, the attacker has full access to all information about the victim model.
+Methods such as PGD attack [34] and AutoAttack [17] generate adversarial examples by leveraging the gradient direction.
+Although this scenario is often unrealistic in practical settings, research in this area is valuable for developing more robust models in the future.
+In contrast, black-box attacks assess the risks of adversarial attacks in real-world scenarios where attackers have limited access to the model’s internals.
+Square Attack [2] and ZO-NGD [54] demonstrate that adversarial examples can be generated by querying the model’s output predictions.
+Additionally, adversarial examples can be crafted through transferability, where models with similar architectures are used to generate adversarial examples that remain effective across different models [41, 13].
 2.2 Defensive Strategies
-Adversarial training is a defensive strategy that aims to find optimal weights against adversarial attacks [8, 33, 34].
+Adversarial training is a defensive strategy that aims to find optimal weights against adversarial attacks.
 It achieves this by generating adversarial examples on the fly during the training phase and optimizing the model’s weights to minimize the losses caused by these examples.
@@
 3 Methodology
-3.1 Motivation
-In this paper, we approach robustness from a theoretical perspective, aiming to demonstrate that all risks posed by adversarial examples are limited while minimizing the additional costs associated with improving robustness.
-Our evaluation is conducted under the white-box scenario, where the target model is capable of defending against various types of known adversarial attacks, including white-box attacks [8, 18, 12], black-box attacks [43], and transfer attacks [44, 45].
-Additionally, we conduct a set of experiments to verify that gradient masking [14] does not occur in our method and to ensure that robustness is not overestimated.
+The main goal of this paper is to enhance the robustness of adversarially trained models by leveraging Lipschitz continuity theory to remap the input domain.
+We propose a data-driven approach that aachieves state-of-the-art performance with minimal extra computational cost, effectively balancing model resilience and training efficiency.
 3.1 Lipschitz Continuity
@@
 Under the $L_{2}$ norm, we have
-$\frac{||f(x^{\text{adv}})-f(x)||_{2}}{||x^{\text{adv}}-x||_{2}}=\frac{(||Wx^{\text{adv}}+b)-(Wx+b)||_{2}}{||\delta||_{2}}=\frac{||W\delta||_{2}}{||\delta||_{2}},$ (4)
+[…]
+[…]
 where $W$ is the weight matrix; and $\delta$ is the distance between $x^{\text{adv}}$ and $x$.
 Therefore, the original optimization problem of minimizing Lipschitz constant is transformed into the following minimization problem:
-$\min_{W}\max_{\delta\neq 0,\delta\in\mathbb{R}^{m}}\frac{||W\delta||_{2}}{||\delta||_{2}}=\min_{W}\sigma_{\text{max}}(W),$ (5)
+[…]
 where $\sigma_{\text{max}}(W)$ represents the largest singular value of the matrix $W$. Notably, there is a relation to eigenvalues:
@@
 In other words, minimizing $\lambda_{\text{max}}(WW^{\dagger})$, the largest eigenvalue of the matrices, can achieve the same objective.
-Rather that minimizing the objective directly, Gershgorin circle theorem provides an alternative solution to estimate the robustness of the given linear system.
+3.2 Forged Function
+One approach to improving the robustness of a neural network is to minimize the objective function defined in (5) to obtain an optimal weight matrix.
+However, we argue that the largest singular value provides only a loose bound for the Lipschitz constant, as adversarial examples typically lie within an $\epsilon$-ball with a small radius, rather than spanning the entire input space.
+This observation suggests an alternative strategy: remapping the input domain to a constrained range, which effectively shrinks the reachable space.
+The restricted set should suppress values in specific dimensions most affected by adversarial examples, thereby minimizing the distribution mismatch between natural and adversarial data, all while maintaining accuracy.
+Therefore, the problem can be decomposed into two parts: (a) identifying an appropriate constrained range, and (b) transforming it into a corresponding weight matrix that ensures a bounded Lipschitz constant.
+We propos a forged function for each layer defined as follows:
+$f^{\text{forge}}_{i}(x)=\begin{cases}xf^{\text{masked}}_{i}(x)\quad&\text{if}\quad|x|\leq c^{\text{th}}_{i},\\ x\quad&\text{otherwise},\end{cases}$ (7)
+where $c^{\text{th}}_{i}$ is a threshold for the $i$-th layer; $f^{\text{masked}}_{i}(x)$ is a scaling function with values in the range $[0,1]$.
+Compared with the original input, the range of the forged function is suppressed if its value is less than the threshold.
+When $c^{\text{th}}_{i}$ is set to 0, the forged function degrades into the original function.
+The simplest definition of scaling function is a step function:
+$f^{\text{step}}(x)=\begin{cases}0\quad&|x|\leq c^{\text{th}}_{i},\\ 1\quad&\text{otherwise}.\end{cases}$ (8)
+However, discontinuous functions might cause unreliable gradients, potentially exposing the defense to BPDA attacks [3].
+Therefore, we propose a smooth function that gradually reduces the value to 0, defined as:
+$f^{\text{smooth}}=\frac{1}{1+\exp(-ax+b)},$ (9)
+where $a$ controls the slope, and $b$ determines the location of the transition zone. Alternatively, a piecewise linear function can simplify this behavior:
+$f^{\text{piecewise}}(x)=\begin{cases}0\quad&|x|\leq dc^{\text{th}}_{i},\\ \frac{|x|-dc^{\text{th}}_{i}}{(1-d)c^{\text{th}}_{i}}\quad&dc^{\text{th}}_{i}\leq|x|\leq c^{\text{th}}_{i},\\ 1\quad&\text{otherwise},\end{cases}$ (10)
+where $d$ is within the range $[0,1]$.
+In (10), the transition zone is represented by a linear function, and as a result, the composed $f^{\text{forge}}$ in (7) becomes a quadratic function within this zone.
+(a) ConvNet
+(b) Transformer
+Figure 1: Insertion points of the forged function. In ConvNets, it is inserted into the residual blocks, while in Transformers, it is inserted into the MLP layers.
+Figure 1 provides a visual representation of potential insertion points for the forged function, while maintaining the integrity of other layers.
+For the ResNet architecture, the forged function is placed before the convolutional layers in each residual block.
+Similarly, for vision transformer architectures, the structure of MLP layers is adapted to seamlessly integrate the forged function.
+However, although the Query (Q), Key (K), and Value (V) computations in the attention layers involve linear projections, we do not apply the forged function to these components due to the more complex implementation involved.
+To understand how the forged function influences the Lipschitz constant, we begin by recalling the Gershgorin Circle Theorem:
 Theorem 1.
@@
 For any layer which can be represented by a linear system, such as convolutional or fully connected layers, robustness can be improved by shrinking the radius of the disk with the largest eigenvalue.
-3.3 Forged Function
-(a) ConvNet
-(b) Transformer
-Figure 2: Insertion points of the forged function. In ConvNets, it is inserted into the residual blocks, while in Transformers, it is inserted into the MLP layers.
-We argue that the largest singular value provides a loose bound for the Lipschitz constant.
-To precisely reflect the robustness of the corresponding observed data, we define the empirical Lipschitz constant that eliminates the influence of space that is never drawn from real data.
-Definition 1.
-Empirical Lipschitz constant:
-$\max_{\delta\neq 0,x\in\mathcal{S}}\frac{||Wx||_{2}}{||x||_{2}}\quad\forall x\in\mathcal{S},$ (8)
-where $\mathcal{S}$ is an observed dataset. As can be seen, the empirical Lipschitz constant on the finite dataset is less than or equal to its Lipschitz constant derived from the theorem.
-Based on Definition 1, we can build robust models by manipulating the output ranges of individual layers, thereby restricting the input domain of the next layer.
-If input vectors do not align with the direction of the eigenvector with the largest eigenvalue, the empirical constant should be bounded.
-Therefore, we proposed a forged function defined as follows:
-$f^{\text{forge}}(x)=\begin{cases}0\quad&\text{if}\quad|x|\leq c^{\text{th}}_{i},\\ x\quad&\text{otherwise},\end{cases}$ (9)
-where $c^{\text{th}}_{i}$ is a threshold for the $i$-th layer.
-Compared with the original functions, the range of the forged function is suppressed if its value is less than the threshold.
-When $c^{\text{th}}_{i}$ is set to 00, the forged function degrades into the original function.
-The forged function aim to reduce the empirical Lipschitz constant of the layers that can be represented as linear systems by remapping the input domain of these layers into a constrained set.
-Figure 2 provides a visual representation of potential insertion points for the forged function, while maintaining the integrity of other layers.
-For the ResNet architecture, the forged function is placed before the convolutional layers in each residual block.
-Similarly, for vision transformer architectures, the structure of MLP layers is adapted to seamlessly integrate the forged function.
-Here is the proof that the largest eigenvalue can be shrunk by the forged function.
-Let $W$ be the weight of the target layer, which can be represented by an $m\times n$ matrix, and t be the input vector.
+Now, let $W$ be the weight of the target layer, which can be represented by an $m\times n$ matrix, and t be the input vector.
 Without loss of generality, we assume that $A=W^{\dagger}W$ and $f^{\text{forge}}(\textbf{t})$ is defined as:
-$f^{\text{forge}}(t_{i})=\begin{cases}0\quad i\leq k\\ t_{i}\quad\text{otherwise},\end{cases}$ (10)
-where $t_{i}$ is the $i$-th element of t and $k$ is a positive number.
+$f^{\text{forge}}(t_{i})=\begin{cases}0\quad i\leq h\\ t_{i}\quad\text{otherwise},\end{cases}$ (12)
+where $t_{i}$ is the $i$-th element of the vector t and $h$ is a positive number.
 Lemma 2.
@@
 Proof.
-Since the first $k$ entries of the vector t are replaced with zeros, above condition can be achieved by replacing the corresponding column vectors of the matrix $A$ with zero vectors.
+Since the first $h$ entries of the vector t are replaced with zeros, above condition can be achieved by replacing the corresponding column vectors of the matrix $A$ with zero vectors.
 Therefore, the entries of $A^{\prime}$ are formulated as
-$a^{\prime}_{ij}=\begin{cases}0\quad j\leq k\\ a_{ij}\quad\text{otherwise}.\end{cases}$ (12)
+$a^{\prime}_{ij}=\begin{cases}0\quad j\leq h\\ a_{ij}\quad\text{otherwise}.\end{cases}$ (14)
 The matrix $A$ is a positive semidefinite matrix, implying that the diagonal entries are non-negative.
-Moreover, with the entry representation of $A^{\prime}$ in (12), we observe that modifications are only applied to the first $k$ columns, while the rest remain unchanged.
-Combining the Gershgorin Circle Theorem, we know that the centers of the first $k$ disks of the matrix $A^{\prime}$ are shifted towards zero.
+Moreover, with the entry representation of $A^{\prime}$ in (14), we observe that modifications are only applied to the first $h$ columns, while the rest remain unchanged.
+Combining the Gershgorin Circle Theorem, we know that the centers of the first $h$ disks of the matrix $A^{\prime}$ are shifted towards zero.
 Additionally, the radii of all disks, the absolute values of the off-diagonal entries in $A^{\prime}$, are shrunk.
@@
 ∎
-Notably, the outputs of $f^{\text{forge}}(\textbf{t})$ vary depending on the inputs, resulting in each input having its own $A^{\prime}$.
-The upper bound of the largest eigenvalue of each matrix $A^{\prime}$ must be not greater than the largest eigenvalue of the matrix $A$.
-With Lamma 2, a precise upper bound of the largest eigenvalue can be obtained by feeding a set of observed images.
-On the contrary, there might be cases in which solving the minimization problem in (5) leads to the theoretical largest eigenvalue being minimized, but the empirical Lipschitz constant remains unchanged.
-The choice of a proper $c^{\text{th}}_{i}$ is a crucial factor in reducing the largest eigenvalue. In this paper, we propose obtaining the value of $c^{\text{th}}_{i}$ through the following equation:
+It is important to note that the proof presented above is specified for each individual input vector, with each input having its own corresponding matrix $A^{\prime}$ that ensures a bounded Lipschitz constant.
+This input-specific proof allows for precise theoretical guarantees on robustness.
+While the corresponding matrix $A^{\prime}$ is auxiliary and used primarily for mathematical proof purposes, there is no need to explicitly compute this matrix during the inference phase.
+3.3 Data-driven Algorithm
+The choice of a proper $c^{\text{th}}_{i}$ represents a critical balance in our approach.
+A larger threshold more effectively constrains the adversarial space and reduces the Lipschitz constant.
+However, excessive restriction can significantly distort the output distribution, potentially degrading overall accuracy on natural examples.
+In this paper, we propose a data-driven approach for determining optimal value through the following equation:
 $c^{\text{th}}_{i}=c^{r}\max(F_{1\rightarrow i}(x))\quad\forall x\in\mathcal{S},$ (15)
-where $\mathcal{S}$ can include all or a subset of images in the training set, $c^{r}$ is a positive number and $F_{1\rightarrow i}(x)$ represents the output of the $i$-th layer.
+where $\mathcal{S}$ can include all or a subset of images in the training set, $c^{r}$ is a positive number within the range $[0,1]$ and $F_{1\rightarrow i}(x)$ represents the output of the $i$-th layer.
 Specifically, each layer has its own $c^{\text{th}}_{i}$, but they share the same hyper-parameter $c^{r}$.
@@
 By default, the mode is set to inference, and the values of $b$ and $c^{r}$ are zero, respectively.
-As a result, the set $\mathcal{M}$ is empty, and the algorithm is degraded to the identical function.
-It is worth emphasizing that by feeding all images in the set $\mathcal{S}$ once in track mode beforehand, the value of $c^{\text{th}}_{i}$ can be obtained.
-The elements satisfying the constraints are appropriately deactivated during inference.
-Notably, this operation does not necessitate gradient computations and incurs minimal time consumption, typically only a few minutes, even when executed on commonly used GPUs.
-In comparison to adversarial training, this process is nearly cost-free.
-The overall procedure shares many similarities with post-pruning techniques.
-Nevertheless, we posit that the proposed function is very similar to the ReLU function, as it suppresses the output values within a specific range, but the defined range in the forged function is adaptive to the observed dataset.
+As a result, the mask $x^{\text{mask}}$ are set to $1$, and the algorithm is degraded to the identical function.
+Our method offers significant efficiency advantages in practical implementation as the threshold $c^{\text{th}}_{i}$ can be obtained by scanning all data in the set $\mathcal{S}$ once in track mode beforehand.
+This process requires no gradient computations and adds minimal computational overhead—typically just minutes even on consumer-grade GPUs.
+Compared to the substantial costs of adversarial training, our approach is essentially cost-free while providing meaningful robustness improvements.
+Conceptually, the proposed forged function operates analogously to a ReLU function, as it suppresses output values within a specific range, but with a key distinction: the defined range in our forged function is adaptively determined based on the characteristics of the observed dataset rather than being fixed.
+This data-driven adaptation enables our method to enhance robustness while preserving performance on natural examples.
 Algorithm 1 Forged Function
-1:  require: Input x, Mode $m$
+1:require: Input x, Mode $m$, Hyper-parameter $c^{r}$
 2:if $m$ is tracking mode then
-3:     $b=\max(b,\textbf{x})$
+3:  $b=\max(b,|\textbf{x}|)$
 4:else
-5:     =$\mathcal{M}=\left\{x\middle|\text{abs}(x)\leq c^{r}b\right\}$
-6:     for all $s\in\mathcal{M}$ do
-7:        $s=0$
-8:     end for
-9:  end if
-10:  return x
+5:  for all $x_{i}\in\textbf{x}$ do
+6:   if $\text{abs}(x_{i})\leq c^{r}b$ then
+7:    $x^{\text{mask}}=f^{\text{masked}}(x)$
+8:   else
+9:    $x^{\text{mask}}$ = 1
+10:   end if
+11:   $x_{i}\leftarrow x^{\text{mask}}x_{i}$
+12:  end for
+13:end if
+14:return x
 4 Experiments
 4.1 Setup
-We evaluated the performance on CIFAR10, CIFAR100, and ImageNet datasets under the white-box scenario with an $L_{\infty}$ norm.
-To ensure comparability of results, we assessed robustness using AutoAttack [12], For CIFAR10 and CIFAR100 datasets, $\epsilon$ is set to $8/255$, while for the ImageNet dataset, $\epsilon$ is set to $4/255$.
-The model weights are publicly accessible from RobustBench.
-The ablation study involves exploring the selection of the optimal $c^{r}$, the combination of various models trained from different techniques, the verification of gradient masking, and assessments of certified adversarial robustness via randomized smoothing.
-Due to the page limit, the full experimental results of the ablation study are listed in Appendix.
-4.2 White-box Evaluation
-4.2.1 Performance Analysis on CIFAR10 and CIFAR100 Datasets
-Table 1: The results of top-3 competitors on Robustbench.
-(a) CIFAR10 dataset
-# Method accnat accAA
-* [11] + Ours 93.20 71.70
-1 [48] 93.27 71.07
-2 [11] 93.25 70.69
-3 [49] 95.19 69.71
-(b) CIFAR100 dataset
-# Method accnat accAA
-* [11] + Ours 74.97 44.00
-1 [11] 75.22 42.67
-2 [49] 83.08 41.80
-3 [50] 73.85 39.18
-The model used in this study is based on WRN-70-16 architecture with SiLU function while generative data were involved during the training phase.
-The value of $c^{\text{th}}_{i}$ is obtained by feeding all images from the training set without any augmentation, and $c^{r}$ was set to $2^{-8}$ for this experiment.
-Tables 1(a) and 1(b) summarize the top-3 competitors on Robustbench for CIFAR10 and CIFAR100 datasets, respectively, where ##\## represent the rankings, our results are marked by the asterisk (*), accnat and accAA denote the accuracy against clean data and adversairal examples generated by AutoAttack, respectively.
-As can be seen, our method combined with WRN-70-16 with SiLU function gains improvement in robustness by at least 0.9% and achieves the best results on Robustbench for both datasets.
-Nevertheless, standard accuracy (accnat) is decreased.
-Many factors might affect the results.
-For example, the single additional hyper-parameter introduced in this study might not provide sufficient granularity to fit all layers in the target model.
-4.2.2 Performance Analysis on ImageNet Dataset
-Table 2: The results of top-3 competitors for ImageNet dataset on Robustbench.
-# Method Architecutre accnat accAA
-* [51] + Ours Swin-L 78.88 60.04
-1 [51] Swin-L 78.92 59.56
-2 [49] ConvNeXtV2-L + Swin-L 81.48 58.50
-3 [51] ConvNeXt-L 78.02 58.48
-In this experiment, we utilized the Swin [52] model architecture, a variant of transformers.
-However, scanning the approximately 1.2 million training images provided by the ImageNet dataset to determine the value of the hyper-parameter introduced in the forged function defined in (13) might take a long time.
-Alternatively, we randomly selected about 5,000 images as the observed images to determine the value of the hyper-parameter.
+We evaluated our method on CIFAR10, CIFAR100, and ImageNet datasets under white-box attack scenarios using AutoAttack [17] with $L_{\infty}$ norm constraints ($\epsilon=8/255$ for CIFAR10/100 and $\epsilon=4/255$ for ImageNet).
+All baseline models were sourced from the public RobustBench repository for reproducibility.
+To determine threshold values $c^{\text{th}}_{i}$, we processed the complete training sets for CIFAR10 and CIFAR100 without data augmentation.
+For the larger ImageNet dataset, scanning all 1.2 million training images would require substantial time.
+Alternatively, we randomly sampled approximately 5,000 representative images to establish thresholds efficiently.
 Ideally, determining the optimal choice of $c^{r}$ requires conducting an ablation study to explore the relationship between the chosen $c^{r}$ and robust accuracy on a validation set.
@@
 The candidate values are selected in a small range centered around this value.
-Tables 2 lists the top-3 competitors on Robustbench for ImageNet dataset, including ranking, architecture, standard accuracy, and robust accuracy against AutoAttack.
-The experimental results demonstrate that the Swin-L model with GELU combined with our method can obtain improvements in robust accuracy and achieve the best result while standard accuracy has a tiny drop.
-This finding verifies that our method can be applied to both convolutional and fully connected layers.
-4.2.3 Combination with Various Models
-The experiment aims to assess the generability of the proposed function on adversarial trained models with identical architecture but from various training strategies and to evaluate the potential cost reduction of adversarial training.
-We integrated the proposed approach with partial models selected from RobustBench, whose weights are obtained directly from the official without any modifications, and also included a model trained by TRADES [53] as a baseline for CIFAR10 dataset.
-The selected models were trained using different techniques, such as adding perturbations in internal layers, retrieving information using knowledge distillation, reducing inefficient training data, or involving additional images from generated models or another dataset.
-Except for the model used in RST-WAP, which is WRN-28-10, the model architecture we utilized is WRN-34-10 with ReLU, as it is the most popular network on the RobustBench leaderboard [13].
-For the white-box evaluation, the value of $c^{r}$ is set to $2^{-7}$.
-Table 3(a) and 3(b) present standard and robust accuracy of models integrated with our method for CIFAR10 and CIFAR100 dataset, respectively.
-In these tables, the column Original indicates the original results reported by RobustBench, and the column Original+Ours demonstrates the results of the proposed method.
+Since most models are represented in 16 bit format, and the widths of fraction bit for FP16 format defined by IEEE-754 standard and BFloat are 10 and 7 bits, respectively, truncated errors might easily occur when performing addition on two numbers with a magnitude difference of $2^{8}$ or higher.
+On the other hand, when $c^{r}$ is set to $2^{-5}$, all models experience a significant drop in standard accuracy, and there is meaningless in evaluating robustness at this configuration.
+We suggest that the candidates of $c^{r}$ are $2^{-8}$, $2^{-7}$ and $2^{-6}$ and the partial results are presented in Table 3.
+Our experimental investigation included ablation studies examining: optimal scaling coefficient selection, compatibility with various adversarially trained models, verification against gradient masking, and certified robustness through randomized smoothing.
+Complete ablation study results are presented in Appendix A due to space constraints.
+4.2 White-box Evaluation
+Method Original Original+Ours
+accnat accAA accnat accAA
+RST-AWP [45] 88.25 60.04 89.50 62.76
+DefEAT [11] 86.54 57.30 87.40 59.55
+LTD [10] 85.21 56.94 85.98 59.25
+AWP [45] 85.36 56.17 86.19 57.85
+TRADES[51] 85.34 52.86 85.78 53.80
+(a) CIFAR10 dataset
+Method Original Original+Ours
+accnat accAA accnat accAA
+EffAug [1] 68.75 31.85 69.14 32.57
+DKLD [19] 64.08 31.65 64.26 32.58
+DefEAT [11] 64.32 31.13 66.42 32.11
+LTD [10] 64.07 30.59 64.29 31.95
+AWP [45] 60.38 28.86 60.63 29.72
+(b) CIFAR100 dataset
+Table 1: Standard and robust accuracy of models integrated with our method.
+This experiment evaluates the generalizability of the proposed method on adversarially trained models with the same architecture but trained using different strategies.
+Additionally, we assess the potential cost reduction from adversarial training.
+We integrated our approach with models selected from RobustBench, trained using various techniques such as adding perturbations in internal layers, retrieving information using knowledge distillation, reducing inefficient training data, or involving additional images from generated models or another dataset.
+Except for the WRN-28-10 model used in RST-WAP, all models were based on the popular WRN-34-10 with ReLU, the leading architecture on the RobustBench leaderboard [18].
+The value of $c^{r}$ is set to $2^{-7}$.
+Table 1(a) and 1(b) present standard and robust accuracy of models integrated with our method for CIFAR10 and CIFAR100 dataset, respectively.
+In these tables, accnat and accAA refer to accuracy on clean data and adversarial examples generated by AutoAttack, respectively.
+The column Original indicates the original results reported by RobustBench, and the column Original+Ours demonstrates the results of the proposed method.
 As indicated in these tables, for CIFAR10 dataset, the proposed method enhances robust accuracy by more than $2\%$ for RST-AWP, DefEAT, and LTD models, while other models receive approximately $1$ to $1.5\%$ improvement in robustness.
@@
 We believe that the proposed solution is general as it achieves great success in models incorporating different training techniques.
-Another advantage of the proposed method that we would like to highlight is that the cost of our approach can almost be ignored compared to the cost of adversarial training as the cost involves only a single pass scan of a set of images to determine the hyper-parameter $c^{th}$.
-This implies that these models can enhance robustness for free.
-Specifically, LefEAT can achieve a robust accuracy of 57.30% by removing inefficient training data.
-By combining LefEAT model with our approach, a robust accuracy of 59.55% can be achieved, which is comparable to RST-AWP (60.04%).
+Additionally, models integrated with our approach incur significantly lower training costs compared to models with similar robustness.
+For example, combining the LefEAT model with our method achieves a robust accuracy of 59.55% can be achieved, which is comparable to RST-AWP (60.04%).
 However, RST-AWP introduces more images from another dataset, resulting in a higher cost in each epoch.
 Similarly, for the CIFAR100 dataset, DefEAT with our proposed method achieves a robust accuracy of 32.11%, which is better than EffAug (31.85%), which involves more complex data augmentation during the training stage.
-This aligns with the suggestion by DefEAT that some data can be removed without hurting robustness.
-Holistically, we believe that our approach might provide a hint during the late phase of adversarial training to drop inefficient weights, resulting in further cost savings or enhanced resilience.
-Another interesting observation presented in these tables is that standard accuracy is also improved for all models on both CIFAR10 and CIFAR100 datasets.
-Although this phenomenon cannot be explained by Lemma 2, we believe that the output range between ReLU and the proposed functions has a high similarity.
-Consequently, this arrangement can maintain the accuracy of clean data.
-Table 3: Standard and robust accuracy of models integrated with our method.
+This result supports DefEAT’s suggestion that unnecessary data can be eliminated without sacrificing robustness.
+Overall, we believe that our approach offers a valuable strategy for pruning inefficient weights in the later stages of adversarial training, potentially leading to further cost savings or enhanced resilience.
+An interesting observation from the tables is that standard accuracy improves for all models on both CIFAR10 and CIFAR100.
+While this is not directly explained by Lemma 2, we hypothesize that the output ranges of ReLU and our proposed function are similar, helping maintain accuracy on clean data.
+Tables 2(a), 2(b), and 2(c) present the top-3 models on RobustBench for the CIFAR10, CIFAR100, and ImageNet datasets, respectively.
+The rankings are denoted by #\#, with our results marked by an asterisk (*).
+For CIFAR10 and CIFAR100 datasets, we used the WRN-70-16 architecture with SiLU activations, incorporating generative data during training.
+As shown, our approach, when applied to the WRN-70-16 model with SiLU, improves robustness by at least 0.9%, achieving the best results on RobustBench for both datasets.
+However, there is a slight decrease in standard accuracy (accnat).
+Several factors may explain this outcome.
+For instance, the additional hyperparameter introduced in this study might not offer sufficient granularity to optimize all layers of the target model effectively.
+Although the proposed method was originally designed for ReLU activations, these results indicate its applicability to other activation functions.
+Future work could explore the design of more specialized functions to further optimize performance with different activation function.
+On the other hand, For the ImageNet dataset, we employed the Swin model [33], a transformer-based architecture.
+Experimental results show that the Swin-L model with GELU activations, combined with our method, improves robust accuracy and achieves the best result, while standard accuracy sees only a marginal decrease.
+This suggests that our approach is effective across both convolutional and transformer-based architectures.
+# Method Architecture accnat accAA
+* [43] + Ours WRN-70-16 93.20 71.70
+1 [37] WRN-70-16 93.27 71.07
+2 [43] WRN-70-16 93.25 70.69
+3 [5] WRN-70-16 95.19 69.71
 (a) CIFAR10 dataset
-Method Original Original+Ours accnat accAA accnat accAA RST-AWP [54] 88.25 60.04 89.50 62.76 DefEAT [39] 86.54 57.30 87.40 59.55 LTD [55] 85.21 56.94 85.98 59.25 AWP [54] 85.36 56.17 86.19 57.85 TRADES[53] 85.34 52.86 85.78 53.80
+# Method Architecture accnat accAA
+* [43] + Ours WRN-70-16 74.97 44.00
+1 [43] WRN-70-16 75.22 42.67
+2 [5] WRN-70-16 83.08 41.80
+3 [19] WRN-70-16 73.85 39.18
 (b) CIFAR100 dataset
-Method Original Original+Ours accnat accAA accnat accAA EffAug [56] 68.75 31.85 69.14 32.57 DKLD [50] 64.08 31.65 64.26 32.58 DefEAT [39] 64.32 31.13 66.42 32.11 LTD [55] 64.07 30.59 64.29 31.95 AWP [54] 60.38 28.86 60.63 29.72
+# Method Architecutre accnat accAA * [32] + Ours Swin-L 78.88 60.04 1 [32] Swin-L 78.92 59.56 2 [5] ConvNeXtV2-L + Swin-L 81.48 58.50 3 [32] ConvNeXt-L 78.02 58.48
+(c) ImageNet dataset
+Table 2: The results of top-3 competitors on Robustbench.
+Method RobustBench $c^{r}=2^{-8}$ $c^{r}=2^{-7}$ $c^{r}=2^{-6}$
+accnat accAA accnat accAA accnat accAA accnat accAA
+TRADES [51] 85.34 52.86 85.57 52.97 85.78 53.80 85.49 55.37
+Table 3: Ablation study of selecting optimal $c^{r}$ for CIFAR10 dataset.
+4.3 Cost Analysis
+For the CIFAR-10 and CIFAR-100 datasets, the total training time for AWP with WRN-34-10 on a V100 GPU is approximately 20 hours.
+In contrast, other methods use larger models and incorporate additional data [43, 37, 5], leading to a total training time of over 200 hours.
+In comparison, scanning the entire CIFAR-10 or CIFAR-100 dataset, or partial training set on ImageNet dataset, takes less than 5 minutes on a V100 GPU.
+Regarding the costs of hyperparameter search, as discussed in Appendix A.1, there are only three candidate parameters to evaluate performance.
+We can efficiently assess white-box performance using partial data from the training set, which significantly reduces the computational time.
+In practice, the total time for hyperparameter search is about 1-3 hours, depending on the model size and the dataset used.
+This demonstrates that our approach incurs significantly lower computational overhead.
 4.3.1 Gradient Masking Verification
@@
     Certified robustness that conducted by random smoothing [16].
-The first item has been examined by AutoAttack as it involves three white-box attacks and one black-box attack.
-By compared robust accuracy shown in Table 1(a), 1(b) and 2, the models combined with the proposed method perform better robust accuracy than the original models.
-It indicates black-box cannot produce more adversarial examples.
-The full experimental results for the rest of experiments can be found in the appendix.
+AutoAttack has examined the first item, which involves three white-box attacks and one black-box attack.
+By comparing robust accuracy shown in Table 2(a), 2(b) and 2(c), the models combined with the proposed method perform better robust accuracy than the original models.
+It indicates black-box attacks cannot produce more adversarial examples.
+The full experimental results for the rest of the experiments can be found in Appendix B.
 The results demonstrate that the proposed algorithm does not violate any of the above rules and the certified robustness improves by our method across most settings.
 From the evidence, we believe that the proposed method does not encounter the gradient masking problem among different hyper-parameters and various models on CIFAR10 and CIFAR100 datasets.
+4.4 Discussion
+Comparison with Lipschitz-constrained methods A common misconception is that minimizing the Lipschitz constant necessarily improves model robustness.
+However, in extreme cases — for example, replacing the weights of a linear layer with an identity matrix — the Lipschitz constant becomes 1, but the resulting model may fail to generalize, yielding zero natural accuracy and thus no meaningful robustness to assess.
+While certified training [35] and Lipschitz-constrained methods [56] can reduce the Lipschitz constant, they often introduce additional regularization terms or complex objectives, making it difficult to balance robustness and performance.
+In contrast, our approach enhances robustness by shrinking the space of adversarial examples, without requiring explicit minimization as in (5).
+Effective adversarial training The proposed function can be constructed in a single pass over the training data, without generating adversarial examples.
+By suppressing values in regions associated with adversarial perturbations, our method reduces the distributional mismatch between clean and adversarial data.
+Although similar ideas have been explored in prior work [4], such methods typically rely on explicit adversarial training.
+Additionally, the empirical results show that when applied to models trained without additional generative data, our method achieves robustness comparable to or better than models that use extensive supplementary data.
+This implicitly suggests that unnecessary data can be eliminated without without degrading prediction quality which is aligned with previous work [11].
+Structured pruning The overall procedure of the proposed algorithm shares many similarities with pruning techniques [25].
+However, we would like to emphasize that the proposals are distinctly different.
+Pruning primarily aims to create a highly sparse model that accelerates inference times or reduces the model size for deployment on edge devices, often without considering robustness.
+In particular, when the pruned model exhibits extremely high sparsity without applying re-training or fine-tuning, natural accuracy drops significantly, implicitly indicating that condition (13) does not hold and that Lemma 2 is not applicable in this case.
+On the other hand, when the pruned model has low sparsity, only a small proportion of weights with values close to zero are eliminated, resulting in insignificant adjustments to the center and radius of the disk.
+The proposed algorithm, on the other hand, can identify crucial elements with minimal cost.
+We believe that investigating the impact of various pruning techniques, such as iterative or post-training methods, on robustness, or combining these techniques with our proposed approach, represents a valuable direction for future research.
 5 Conclusion
-In this paper, we recap how robustness is certified by the theorem of Lipschitz continuity.
-We introduce the concept of the empirical Lipschitz constant, which minimizes the influence of the space not drawn from real data, resulting in a precise estimation of the robustness of the corresponding observed data.
-We prove that by remapping the input domain of a specific layer to a constrained range, the Lipschitz constant can be shrunk, leading to better robustness.
-The proposed function introduces only one parameter, the value of which can be determined by scanning the training data once, without re-training or fine-tuning.
-Compared with adversarial training, the proposed method is almost cost-free.
-The experimental results suggest that our method can be combined with various existing methods and achieve robustness improvements, and no gradient masking occurs in our algorithm.
-Furthermore, our method can achieve the best robust accuracy for CIFAR10, CIFAR100, and ImageNet datasets on the RobustBench leaderboard.
+In this paper, we revisited robustness certification through the concept of Lipschitz continuity.
+While existing Lipschitz-constrained methods reduce the Lipschitz constant using regularization or complex objectives, they often struggle to strike a balance between robustness and performance.
+In contrast, from a theoretical perspective, we demonstrated that remapping inputs to a constrained set can mitigate the influence of regions outside the real data distribution, leading to improved robustness.
+Our method is data-driven and highly efficient, requiring only a single pass through the dataset to determine the appropriate parameters for constructing the forged function without fine-tuning or gradient computations.
+This procedure is computationally efficient and almost cost-free.
+The experimental results highlight that our method can be seamlessly integrated with various existing techniques to enhance robustness.
+When combined with models trained without additional generative data, our method achieves robustness comparable to or exceeding that of models using extensive supplementary data.
+These results open a promising direction for significantly reducing computational costs while maintaining or improving defensive capabilities of robust neural networks.
 Numerous future directions merit exploration.
-Firstly, due to the property of maximization, the proposed function might easily be influenced by outliers.
-Designing a better function is an interesting research topic.
-Secondly, exploring the combination with various activation functions, different model architectures or large-scale datasets would be beneficial.
-Lastly, it is worth investigating to understand the theoretical reasons why our proposed function improves standard accuracy.
+For instance, investigating the integration of the proposed remapping function with different activation functions, model architectures, or large-scale datasets would be beneficial.
+Additionally, it would be worthwhile to explore the theoretical foundations behind the observed improvements in standard accuracy introduced by our method.
+Acknowledgments
+We thank the founding support from National Science and Technology Council ( 113-2622-E-007 -018 ) and the computing infrastructure provided by the National Center for High-Performance Computing, National Institutes of Applied Research (NIAR), Taiwan.
 Appendix A Ablation Study
@@
 The results on CIFAR10 and CIFAR100 are presented in Table 4 and Table 5, respectively.
-Moreover, the results of accuracy against CW attack on $L_{\infty}$ norm for CIFAR10 and CIFAR100 datasets are presented in Tables 6(a) and 6(b), respectively.
+Moreover, the results of accuracy against CW attack on $L_{\infty}$ norm for CIFAR10 and CIFAR100 datasets are presented in Tables 7(a) and 7(b), respectively.
 As can be seen, when $c^{r}$ is set to $2^{-8}$, all models achieve better standard accuracy and robust accuracy.
@@
 Besides, the hyper-parameter is determined by choosing the maximum value appearing in the dataset.
-Table 4: Ablation study of selecting optimal $c^{r}$ for CIFAR10 dataset.
 Method RobustBench $c^{r}=2^{-8}$ $c^{r}=2^{-7}$ $c^{r}=2^{-6}$
@@
 TRADES 85.34 52.86 85.57 52.97 85.78 53.80 85.49 55.37
-Table 5: Ablation study of selecting optimal $c^{r}$ for CIFAR100 dataset.
+Table 4: Ablation study on selecting optimal $c^{r}$ for CIFAR10 dataset.
 Method RobustBench $c^{r}=2^{-8}$ $c^{r}=2^{-7}$ $c^{r}=2^{-6}$
@@
 AWP 60.38 28.86 60.18 29.10 60.63 29.72 60.71 30.82
+Table 5: Ablation study on selecting optimal $c^{r}$ for CIFAR100 dataset.
+A.2 Scaling Training Data for Hyperparameter Tuning
+This experiment investigates how the number of training images used for hyperparameter selection affects both standard accuracy and robust accuracy.
+As shown in Table 6, robust accuracy decreases slightly as the amount of training data increases, but it consistently remains higher than that of the competing methods.
+# Method Training Data [#] accnat accAA 1 [32] + Ours 5,000 78.88 60.04 2 [32] + Ours 10,000 78.92 59.97 3 [32] + Ours 20,000 78.90 59.81 4 [32] - 78.92 59.56 5 [5] - 81.48 58.50
+Table 6: Ablation study on the impact of scaling training data for ImageNet dataset.
 Appendix B Full Experimental Results of Gradient Masking Verification
-Table 6: The robust accuracy against CW attack on $L_{\infty}$ norm.
-(a) CIFAR10 dataset
 Method Origin $c^{r}$
@@
 TRADES 56.10 56.52 58.18 63.62
-(b) CIFAR100 dataset
+(a) CIFAR10 dataset
 Method Origin $c^{r}$
@@
 AWP 34.56 35.20 35.94 40.40
-Table 7(a) and 7(b) present the robust accuracy against adversarial examples generated by the original models on CIFAR10 and CIFAR100 datasets, respectively.
+(b) CIFAR100 dataset
+Table 7: The robust accuracy against CW attack on $L_{\infty}$ norm.
+Table 9(a) and 9(b) present the robust accuracy against adversarial examples generated by the original models on CIFAR10 and CIFAR100 datasets, respectively.
 As observed, none of the models showed lower robust accuracy than the original model.
 It indicates that adversarial examples can be efficiently crafted by utilizing the gradients from the victim models.
-Table 8 and 9 presents the robust accuracy against FGSM and PGD attacks among different radii of the $\epsilon$-ball on the CIFAR10 and CIFAR100 datasets, respectively.
+Table 10 and 11 presents the robust accuracy against FGSM and PGD attacks among different radii of the $\epsilon$-ball on the CIFAR10 and CIFAR100 datasets, respectively.
 As observed, the robust accuracy against FGSM, a one-step attack, is always higher than the robust accuracy against PGD, an iterative attack.
@@
 As the input distribution drawn from random smoothing and the observed data might have discrepancies, this could result in fluctuations in robustness.
+Alternatively, we can empirically assess the gradient masking using Lemma 2.
+The first step is the same as usual: obtaining the hyperparameter $c^{\text{th}_{i}}$ for each layer.
+The second step involves counting the occurrences of specific elements in the input vectors that are remapped to 0.
+The corresponding columns of the weight matrix are then replaced with zero vectors.
+Table 8 presents the results, where [32] + Ours + Weight refers to the alternative implementation.
+As shown, the alternative implementation achieves better robust accuracy compared to the original model.
+However, the proposed method outperforms the alternative implementation.
+This is because the proposed method offers the flexibility to remap the input vectors on a sample-wise granularity, while the alternative implementation replaces the corresponding columns of the weight matrix, which has a broader influence on all samples.
+Method Architecutre accnat accAA [32] + Ours Swin-L 78.88 60.04 [32] + Ours + Weight Swin-L 78.82 59.71 [32] Swin-L 78.92 59.56
+Table 8: Gradient masking verification by the weight adjustment.
 (a) RST-AWP
@@
 Figure 2: Certified robustness that conducted by random smoothing.
-Table 7: The robust accuracy against adversarial examples generated by the original models.
-(a) CIFAR10 dataset
 Method Origin $c^{r}$
@@
 TRADES 52.86 55.55 55.09 58.68
-(b) CIFAR100 dataset
+(a) CIFAR10 dataset
 Method Origin $c^{r}$
@@
 AWP 28.86 29.88 32.18 36.67
-Table 8: The robust accuracy against FGSM and PGD attacks among different radii of $\epsilon$-ball on CIFAR10 dataset.
+(b) CIFAR100 dataset
+Table 9: The robust accuracy against adversarial examples generated by the original models.
 Method $c^{r}$ Attack $\epsilon$
@@
 PGD 83.23 81.33 75.81 64.56 36.76 3.37 0.02 0.0
-Table 9: The robust accuracy against FGSM and PGD attacks among different radii of $\epsilon$-ball on CIFAR100 dataset.
+Table 10: The robust accuracy against FGSM and PGD attacks among different radii of $\epsilon$-ball on CIFAR10 dataset.
 Method $c^{r}$ Attack $\epsilon$
@@
 PGD 58.19 55.66 50.91 42.37 25.53 5.68 0.09 0.0
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The main claim is stated in the abstract and highlighted in the introduction.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: We leave some topics unexplored for further work as described in the conclusion.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: The proof of Lemma 2 is shown in Section 2.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: The experiment were conducted by the public leaderboard RobusBench. The process is standardized and hardware-independent. The results can be fare compared to the competitiers.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [N/A]
-  24.
-Justification: The model architecture and pre-trained weights proposed in this work will be submitted to Robusbench.
-This information will be published once the results are certified by Robusbench.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: The assessments utilized by RobustBench are standardized.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [N/A]
-  34. Justification: The assessments utilized by RobustBench are standardized. The impact of Statistical variances have been included in the assessments.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [N/A]
-  39.
-Justification: The assessments utilized by RobustBench are hardware-independent.
-Experiments can be performed on any GPU, but powerful GPUs can complete the evaluation in shorter execution times.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: This work follows the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A]
-  49. Justification: The main goal of this work is to design reliable models. This work is unlikely to have a negative social impact.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: This models used in this works are existing public models. Therefore, this work poses no such risks.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: The related works and RobustBench are correctly cited in this work.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64. Justification: This work does not release new assets.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: This work does not involve crowdsourcing nor research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: This work does not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+Table 11: The robust accuracy against FGSM and PGD attacks among different radii of $\epsilon$-ball on CIFAR100 dataset.
+Appendix C Quantitative Analysis of Empirical Lipschitz Constant
+(a) FC1 layers
+(b) FC2 layers
+Figure 3: The average proportion of pruned activations depends on the location of each linear layer. FC1 and FC2 refer to the first and second fully connected layers in MLP blocks, respectively.
+Our proposed approach, however, is data-driven, relying on observed data to automatically determine the appropriate parameters for constructing the forged function.
+The sparsity of the forged vectors is a crucial factor influencing the magnitude of the Lipschitz constant for the corresponding layers, although it is not the only factor.
+Figure 3 illustrates the average proportion of pruned activations, which varies depending on the location of each linear layer.
+FC1 and FC2 refer to the first and second fully connected layers in the MLP blocks, respectively.
+As shown, the proportion of pruned activations is approximately 20% for the FC1 layers, except for the last layer.
+This suggests that the forged function alone cannot significantly reduce the magnitude of the Lipschitz constant.
+On the other hand, the pruned rates for the FC2 layer range from 30% to 95%, depending on the layer’s location.
+However, we emphasize that a higher pruning rate does not necessarily lead to a substantial reduction in the empirical Lipschitz constant.
+This is because, although the Lipschitz constant estimated from most of the data may decline, the empirical Lipschitz constant is based on the worst-case scenario across the entire observed dataset.
+Experimental results show that, after applying our method, the eigenvalue of the worst-case scenario is approximately 95% of the original eigenvalue.